OpenStudy (anonymous):

solve the system by the method of substitution -3x-3y+2z=-7 z=1 -2x-3y+z=-6 Options A-(2,1,-1) B-(2,-1,1) C-(-2,1,1) D-(2,1,1) PLEASE SHOW WORK????? PLEASE HELP ME????

hero (hero):

Hint: replace z with 1 in the two given equations.

OpenStudy (anonymous):

substitute z by 1 on equation 1 and equation 3.....then it will be simultaneous equatuon which is grade 10 staff

OpenStudy (anonymous):

-3x-3y+2z=-7.....(i) z=1......(ii) -2x-3y+z=-6.....(iii) substitute z=1 in equation (i) and (iii) -3x-3y+2=-7.....(i') -2x-3y+1=-6.....(iii') they result to: -3x-3y=-9.....(i') -2x-3y=-7.....(iii') from (i') -3y= -9+3x so substituting in (iii') -2x-9+3x = -7 x =2 remember: -3y= -9+3x so 3y=9-3x 3y = 9-6 3y = 3 y=1 so solution: (x,y,z)= (2,1,1), that's substitution, follow?